# Adaptive Sampling for Hydrodynamic Stability

### v1

@@ Abstract
An adaptive sampling approach for efficient detection of bifurcation boundaries in parametri-zed fluid flow problems is presented herein. The study extends the machine-learning approach of Silvester (Machine Learning for Hydrodynamic Stability, arXiv:2407.09572), where a classifier network was trained on preselected simulation data to identify bifurcated and nonbifurcated flow regimes. In contrast, the proposed methodology introduces adaptivity through a flow-based deep generative model that automatically refines the sampling of the parameter space. The strategy has two components: a classifier network maps the flow parameters to a bifurcation probability, and a probability density estimation technique (KRnet) for the generation of new samples at each adaptive step. The classifier output provides a probabilistic measure of flow stability, and the Shannon entropy of these predictions is employed as an uncertainty indicator. KRnet is trained to approximate a probability density function that concentrates sampling in regions of high entropy, thereby directing computational effort towards the evolving bifurcation boundary. This coupling between classification and generative modeling establishes a feedback-driven adaptive learning process analogous to error-indicator based refinement in contemporary partial differential equation solution strategies. Starting from a uniform parameter distribution, the new approach achieves accurate bifurcation boundary identification with significantly fewer Navier–Stokes simulations, providing a scalable foundation for high-dimensional stability analysis.
@@ 1 Introduction
The study of hydrodynamic stability has a long history, forming one of the central pillars of fluid mechanics and applied mathematics (see [joseph2013stability, landau1987fluid, MR604359, MR1801992]). Classical linear stability theory studies the evolution of small perturbations to a steady base flow and determines whether such disturbances decay or amplify over time (see [MR604359, MR1801992, MR1883627]). When perturbations grow, their amplification can be indicative of a loss of stability, often manifested as a bifurcation marking the transition away from the underlying steady or symmetric state. This theory proceeds by linearising the governing Navier–Stokes equations about a reference solution and studying an associated eigenvalue problem, which provides a computational framework for tracking the critical eigenvalues that mark the onset of instability (see [MR1883627, Wubs2023]). However, linear theory provides information only about the response to infinitesimal disturbances and therefore cannot quantify the role of finite-amplitude perturbations in the promotion of instability. Specifically, eigenvalue analysis offers only local insight into stability and does not account for the transient growth of finite perturbations due to non-normality of the associated eigenfunctions. These limitations motivate the exploration of data-driven approaches for illustrating flow stability and mapping bifurcation behaviour.
@@
A growing body of work has explored machine-learning methods for analysing dynamical systems and detecting bifurcation behaviour. Several recent studies have shown that data-driven models can efficiently identify multiple bifurcations in low-dimensional systems [szep2021parameter, mogharabin2025bifurcation]. Related approaches have employed machine-learning techniques to classify trajectories as regular or chaotic from time-series data, and to approximate Lyapunov exponents using neural-network-based surrogates [celletti2022classification, MR4844871]. In the context of fluid mechanics, physics-informed neural networks have attracted significant attention following the pioneering work of Raissi et al. [MR4265157], although subsequent studies have highlighted the challenges of applying PINNs to the Navier–Stokes equations [wang2023expert]. Other machine-learning strategies for steady-state flow bifurcations rely on reduced-order modelling approaches [MR4547235], which require extensive snapshot datasets and introduce truncation errors that are difficult to quantify.
@@
More recently, the results in [silvester2024machine] demonstrate that even relatively simple neural network classifiers can accurately distinguish between bifurcated and nonbifurcated flow states when trained on sufficiently informative data, supporting the view that supervised learning can serve as an effective alternative to classical eigenvalue-based stability analysis. One issue, however, is that the training sets in this work were assembled manually using dense parameter sweeps informed by prior knowledge of where the transition was expected to occur. Since each labelled sample required a full computational fluid dynamics (CFD) simulation the generation of a reliable dataset demanded a substantial number of computationally expensive flow solves. This limitation motivates the development of sampling strategies that automatically concentrate computational effort in the most informative regions of the parameter domain.
@@
In this work, we address this challenge by extending the supervised bifurcation classification approach to a fully adaptive, data-efficient methodology. Whereas the earlier study relied on a fixed, manually assembled training set, the present work introduces an adaptive sampling mechanism that dynamically concentrates new simulations where they are most informative. Our approach couples a neural-network classifier with a normalising-flow-based generative model (KRnet) [tang2020deep, MR4385873] that identifies regions of high predictive uncertainty and selectively proposes new parameter points where additional CFD simulations are most valuable. This adaptive refinement process progressively enriches the training set around the evolving stability boundary, improving classifier accuracy while substantially reducing the number of high-fidelity simulations required. The resulting method automates the exploration of nonlinear bifurcation structures in multidimensional parameter domains without relying on eigenvalue computations, or dense parameter sweeps.
@@
In all problems considered in this study, the labelled flow states used to train the classifier are obtained from high-fidelity simulations performed with the IFISS software package [silvester2023ifiss]. Each solution is categorised as bifurcated or nonbifurcated using the kinetic-energy criterion introduced in [silvester2024machine], which has been shown to provide a robust and consistent labelling strategy. This ensures that the adaptive sampling procedure operates on physically meaningful classifications.
@@ 2.1 Symmetry-breaking Channel Flow
When $\delta=0$, the flow remains symmetric until the critical Reynolds number is reached. For nonzero $\delta$, the perturbation biases the flow toward one of the asymmetric steady states. Representative symmetric and asymmetric steady-state flow fields for this configuration can be found in [silvester2024machine].
@@ 2.2 Rayleigh–Bénard Convection Problem
When the imposed temperature difference is sufficiently small, thermal conduction dominates and the flow remains motionless. As the temperature increases, the conductive equilibrium eventually loses stability, giving rise to two possible steady convection states corresponding to clockwise and counterclockwise circulation. These two states form a pitchfork bifurcation pair, and representative flow fields illustrating each branch can be found in [silvester2024machine].
@@ 2.3 A Differentially Heated Cavity Problem
This problem concerns natural convection in a differentially heated rectangular cavity, a configuration that undergoes a Hopf bifurcation from a steady flow to a time-periodic oscillatory regime at sufficiently large Rayleigh numbers. The physical motivation for this problem originates from the classical laboratory experiments of Elder [elder1965laminar], who examined convection in a tall, narrow cavity filled with a highly viscous fluid. The working fluid in those experiments had an effective Prandtl number close to $\mathrm{Pr}\approx 1000$, characteristic of glycerol.
The computational setup considered here follows Elder’s geometry: a cavity of aspect ratio 0.051:10.051{:}1, with the left vertical boundary ($x=0$) held at a higher temperature and the right boundary ($x=0.051$) cooled, while the horizontal walls remain thermally insulated. At low Rayleigh numbers the flow settles into a steady clockwise primary recirculation pattern. As the Rayleigh number increases past a critical value, the steady state loses stability through a Hopf bifurcation and the flow transitions to a time-periodic oscillatory motion. For $\mathrm{Pr}=1000$, this critical transition is observed numerically near $\mathrm{Ra}\approx 2.88\times 10^{9}$ [Wubs2023, p. 291]. To investigate a broader class of fluids ranging from water to glycerol, the bifurcation structure is analyzed in the two-parameter space $(\mathrm{Ra},\mathrm{Pr})$ with $\mathrm{Pr}\in[7.1,1000]$.
@@ 3 Methodology
We extend the machine-learning-based bifurcation classification strategy proposed in [silvester2024machine] into an adaptive and data-efficient approach. While the previous approach employed a neural network trained on a fixed dataset generated by carefully selecting parameter values in regions of the parameter space where transitions from stability to instability were expected, the present work begins with a uniformly sampled set of parameters covering the entire domain. An iterative procedure is then introduced that couples a classification network with a flow-based generative model to guide the sampling of new parameter points.
@@ 3.1 Generation of a Labeled Dataset
As in [silvester2024machine], the learning scheme is based on a two-stage methodology. In the first stage, approximations of the flow solutions are computed by discretising the governing Navier–Stokes equations in space using mixed finite elements and in time using adaptive time stepping. In contrast to [silvester2024machine], where parameter vectors $\boldsymbol{\lambda}=[\lambda_{1},\lambda_{2}]$ were sampled in the neighbourhood of the expected bifurcation boundary, the present study begins with a set of uniformly distributed parameters covering the entire two-dimensional parameter domain. For each parameter combination, the corresponding solution is obtained starting from a quiescent base flow. A classification test is then applied to the computed flow field to determine whether it represents the base (symmetric) flow or whether the flow has undergone a bifurcation. The result is a labelled dataset
@@
where $\ell_{i}\in\{0,1\}$ denotes the label associated with the $i$th parameter combination, indicating nonbifurcated ($\ell_{i}=0$) or bifurcated ($\ell_{i}=1$) behaviour respectively. The labelling procedure follows exactly that described in [silvester2024machine] and is not repeated here. This initial dataset serves as the input to the classifier network described in Section 3.2, which is subsequently refined through adaptive sampling using the generative model introduced in Section 3.3.
@@ 3.2 Classifier Network for Bifurcation Prediction
The second stage of the methodology corresponds to the generation of a surrogate model for the bifurcation structure using a neural network classifier. The network architecture follow exactly those described in [silvester2024machine].
@@
where $\boldsymbol{\lambda}_{\min}$ and $\boldsymbol{\lambda}_{\max}$ denote the lower and upper bounds of the parameter space, respectively. Unlike the normalization strategy adopted in [silvester2024machine], where each parameter component was standardized by subtracting its mean and dividing by its standard deviation, the present work employs a fixed affine scaling to the interval $1,1]^{2}$. This approach ensures consistency with the domain required by the generative model and simplifies uniform numerical treatment of all parameters. The normalized dataset is expressed as
@@ 3.3 Adaptive Sampling with Flow-Based Deep Generative Model
The second component of the proposed approach is a flow-based deep generative model, referred to as KRnet [tang2020deep, MR4385873]. KRnet is a type of normalizing flow that learns an explicit probability density function (PDF) by constructing an invertible mapping between a simple latent space and a complex target space. This invertibility enables both the evaluation of sample likelihoods and the efficient generation of new samples. Compared with adversarial or variational approaches [gulrajani2017improved, goodfellow2014, kingma2013auto], normalizing flow models provide a tractable likelihood formulation, allowing their parameters to be optimized directly through likelihood maximization. However, since the parameter space in the present framework is bounded, the probability distribution to be modeled is supported on a compact domain, whereas flow-based generative models are generally defined over the entire space. To handle the bounded nature of the parameter domain, a smooth bijective mapping is introduced to transform it into an unbounded space where the flow transformation is applied.
@@
Here, the mapping $f$ is realized using the KRnet architecture, which employs a triangular transformation based on the Knothe–Rosenblatt rearrangement [MR2607321]. This design makes KRnet easily scalable to higher-dimensional sampling problems, which is particularly advantageous when the parameter space extends beyond two dimensions.
@@ 3.4 Adaptive Sampling Procedure
For completeness, we briefly summarize the initialization of both models, as this determines how their parameters evolve across adaptive stages. The classifier network is initialized only once at the beginning of the adaptive procedure: all hidden-layer weights and biases are drawn from a zero-mean normal distribution, while the output-layer weights and biases are drawn from a uniform distribution on [1,1][-1,1]. KRnet is likewise initialized only once. The neural networks appearing in the affine coupling layers use Glorot normal initialization [glorot2010understanding] for their weights, and zero initialization for their biases. Once this initialization is performed, the parameters obtained at each stage are retained and used as the starting point for the subsequent adaptive stage.
@@ 4.1 Symmetry-Breaking Channel Flow
For this problem, the initial dataset is generated on a uniform grid in the $(\mathcal{R},\delta)$ parameter space, consisting of $8$ uniformly spaced Reynolds numbers in the range $\mathcal{R}\in[160,230]$ and $16$ logarithmically spaced inlet perturbations $\delta\in[10^{-5},10^{-2}]$, providing $128$ labeled samples. These labels are obtained from Navier–Stokes solutions computed using the IFISS software and classified according to the symmetry criterion of [silvester2024machine]. The classifier neural network used in this example has a single hidden layer with $32$ neurons and a softmax output layer, trained for $n_{e}^{(c)}=3000$ epochs using stochastic gradient descent (batch size $1$, learning rate $0.15$). To construct the initial training set for KRnet, a set of $n_{\text{cand}}=6000$ additional parameter points sampled uniformly in the normalized domain is evaluated by the classifier to compute the entropy weights $w(\boldsymbol{x})$. KRnet employs $6$ affine coupling layers, each equipped with two fully connected subnetworks of $24$ neurons per layer, and is trained using the WNLL loss with the ADAM optimizer (batch size $500$, learning rate $10^{-3}$) for $n_{e}^{(g)}=1000$ epochs. In total, two adaptive refinement steps are carried out, and at each step the KRnet contributes $n_{\text{new}}=50$ additional parameter points to the training set.
@@
A closer inspection of the classification patterns reveals how the adaptive strategy responds to evolving regions of uncertainty. In the initial training dataset, the classifier exhibits its greatest ambiguity near $\mathcal{R}\approx 218$ and $\log_{10}(\delta)\in[-5,-3.5]$, where the bifurcation curve cannot be reliably captured by the coarse grid alone. Consequently, KRnet focuses new samples in this neighbourhood, generating points precisely along the transition curve and producing a sharper and more clear boundary after the first adaptive step. Once this region has been resolved, the remaining uncertainty shifts toward larger perturbations, specifically for $\log_{10}(\delta)\in[-2.75,-2.5]$ and Reynolds numbers $\mathcal{R}\approx 170$$180$. The second adaptive step again directs sampling to this newly uncertain region. These focused refinements collectively highlight the effectiveness of the adaptive sampling approach.
@@
Figure 2 shows the KRnet-generated points at the first two adaptive refinement steps. As discussed thoroughly in the methodology section, KRnet naturally concentrates samples near regions of high classifier uncertainty, which here correspond to the bifurcation boundary. The left plot shows the initial training set consisting of 128 uniformly spaced samples in the $(\mathcal{R},\log_{10}(\delta))$ parameter domain. The middle plot shows the first adaptive update, where KRnet contributes 50 new points, all lying strictly inside the admissible domain. The right plot shows the second adaptive update, containing another 50 KRnet-generated points, of which only a single point falls slightly outside the domain; this point is marked in red. Because such occurrences were rare and had negligible impact on training, no explicit projection step was required, and the single outside-domain point was simply not included in the training set.

### v2

@@ Abstract
An adaptive sampling approach for efficient detection of bifurcation boundaries in parametrized fluid flow problems is presented herein. The study extends the machine-learning approach of Silvester (J. Comput. Phys., 553 (2026), 114743), where a classifier network was trained on preselected simulation data to identify bifurcated and nonbifurcated flow regimes. In contrast, the proposed methodology introduces adaptivity through a flow-based deep generative model that automatically refines the sampling of the parameter space. The strategy has two components: a classifier network maps the flow parameters to a bifurcation probability, and a probability density estimation technique (KRnet) for the generation of new samples at each adaptive step. The classifier output provides a probabilistic measure of flow stability, and the Shannon entropy of these predictions is employed as an uncertainty indicator. KRnet is trained to approximate a probability density function that concentrates sampling in regions of high entropy, thereby directing computational effort towards the evolving bifurcation boundary. This coupling between classification and generative modeling establishes a feedback-driven adaptive learning process analogous to error-indicator based refinement in contemporary partial differential equation solution strategies. Starting from a uniform parameter distribution, the new approach achieves accurate bifurcation boundary identification with significantly fewer Navier–Stokes simulations, providing a scalable foundation for high-dimensional stability analysis.
@@ 1 Introduction
The study of hydrodynamic stability has a long history, forming one of the central pillars of fluid mechanics and applied mathematics (see [9, 11, 4, 16]). Classical linear stability theory studies the evolution of small perturbations to a steady base flow and determines whether such disturbances decay or amplify over time (see [4, 16, 3]). When perturbations grow, their amplification can be indicative of a loss of stability, often manifested as a bifurcation marking the transition away from the underlying steady or symmetric state. This theory proceeds by linearising the governing Navier–Stokes equations about a reference solution and studying an associated eigenvalue problem, which provides a computational framework for tracking the critical eigenvalues that mark the onset of instability (see [3, 23]). However, linear theory provides information only about the response to infinitesimal disturbances and therefore cannot quantify the role of finite-amplitude perturbations in the promotion of instability. Specifically, eigenvalue analysis offers only local insight into stability and does not account for the transient growth of finite perturbations due to non-normality of the associated eigenfunctions. These limitations motivate the exploration of data-driven approaches for illustrating flow stability and mapping bifurcation behaviour.
@@
A growing body of work has explored machine-learning methods for analysing dynamical systems and detecting bifurcation behaviour. Several recent studies have shown that data-driven models can efficiently identify multiple bifurcations in low-dimensional systems [19, 13]. Related approaches have employed machine-learning techniques to classify trajectories as regular or chaotic from time-series data, and to approximate Lyapunov exponents using neural-network-based surrogates [2, 12]. In the context of fluid mechanics, physics-informed neural networks have attracted significant attention following the pioneering work of Raissi et al. [15], although subsequent studies have highlighted the challenges of applying PINNs to the Navier–Stokes equations [22]. Other machine-learning strategies for steady-state flow bifurcations rely on reduced-order modelling approaches [14], which require extensive snapshot datasets and introduce truncation errors that are difficult to quantify.
@@
More recently, the results in [18] demonstrate that even relatively simple neural network classifiers can accurately distinguish between bifurcated and nonbifurcated flow states when trained on sufficiently informative data, supporting the view that supervised learning can serve as an effective alternative to classical eigenvalue-based stability analysis. One issue, however, is that the training sets in this work were assembled manually using dense parameter sweeps informed by prior knowledge of where the transition was expected to occur. Since each labelled sample required a full computational fluid dynamics (CFD) simulation the generation of a reliable dataset demanded a substantial number of computationally expensive flow solves. This limitation motivates the development of sampling strategies that automatically concentrate computational effort in the most informative regions of the parameter domain.
@@
In this work, we address this challenge by extending the supervised bifurcation classification approach to a fully adaptive, data-efficient methodology. Whereas the earlier study relied on a fixed, manually assembled training set, the present work introduces an adaptive sampling mechanism that dynamically concentrates new simulations where they are most informative. Our approach couples a neural-network classifier with a normalising-flow-based generative model (KRnet) [20, 21] that identifies regions of high predictive uncertainty and selectively proposes new parameter points where additional CFD simulations are most valuable. This adaptive refinement process progressively enriches the training set around the evolving stability boundary, improving classifier accuracy while substantially reducing the number of high-fidelity simulations required. The resulting method automates the exploration of nonlinear bifurcation structures in multidimensional parameter domains without relying on eigenvalue computations, or dense parameter sweeps.
@@
In all problems considered in this study, the labelled flow states used to train the classifier are obtained from high-fidelity simulations performed with the IFISS software package [17]. Each solution is categorised as bifurcated or nonbifurcated using the kinetic-energy criterion introduced in [18], which has been shown to provide a robust and consistent labelling strategy. This ensures that the adaptive sampling procedure operates on physically meaningful classifications.
@@ 2.1 Symmetry-breaking Channel Flow
When $\delta=0$, the flow remains symmetric until the critical Reynolds number is reached. For nonzero $\delta$, the perturbation biases the flow toward one of the asymmetric steady states. Representative symmetric and asymmetric steady-state flow fields for this configuration can be found in [18].
@@ 2.2 Rayleigh–Bénard Convection Problem
When the imposed temperature difference is sufficiently small, thermal conduction dominates and the flow remains motionless. As the temperature increases, the conductive equilibrium eventually loses stability, giving rise to two possible steady convection states corresponding to clockwise and counterclockwise circulation. These two states form a pitchfork bifurcation pair, and representative flow fields illustrating each branch can be found in [18].
@@ 2.3 A Differentially Heated Cavity Problem
This problem concerns natural convection in a differentially heated rectangular cavity, a configuration that undergoes a Hopf bifurcation from a steady flow to a time-periodic oscillatory regime at sufficiently large Rayleigh numbers. The physical motivation for this problem originates from the classical laboratory experiments of Elder [5], who examined convection in a tall, narrow cavity filled with a highly viscous fluid. The working fluid in those experiments had an effective Prandtl number close to $\mathrm{Pr}\approx 1000$, characteristic of glycerol.
The computational setup considered here follows Elder’s geometry: a cavity of aspect ratio 0.051:10.051{:}1, with the left vertical boundary ($x=0$) held at a higher temperature and the right boundary ($x=0.051$) cooled, while the horizontal walls remain thermally insulated. At low Rayleigh numbers the flow settles into a steady clockwise primary recirculation pattern. As the Rayleigh number increases past a critical value, the steady state loses stability through a Hopf bifurcation and the flow transitions to a time-periodic oscillatory motion. For $\mathrm{Pr}=1000$, this critical transition is observed numerically near $\mathrm{Ra}\approx 2.88\times 10^{9}$ [23, p. 291]. To investigate a broader class of fluids ranging from water to glycerol, the bifurcation structure is analyzed in the two-parameter space $(\mathrm{Ra},\mathrm{Pr})$ with $\mathrm{Pr}\in[7.1,1000]$.
@@ 3 Methodology
We extend the machine-learning-based bifurcation classification strategy proposed in [18] into an adaptive and data-efficient approach. While the previous approach employed a neural network trained on a fixed dataset generated by carefully selecting parameter values in regions of the parameter space where transitions from stability to instability were expected, the present work begins with a uniformly sampled set of parameters covering the entire domain. An iterative procedure is then introduced that couples a classification network with a flow-based generative model to guide the sampling of new parameter points.
@@ 3.1 Generation of a Labeled Dataset
As in [18], the learning scheme is based on a two-stage methodology. In the first stage, approximations of the flow solutions are computed by discretising the governing Navier–Stokes equations in space using mixed finite elements and in time using adaptive time stepping. In contrast to [18], where parameter vectors $\boldsymbol{\lambda}=[\lambda_{1},\lambda_{2}]$ were sampled in the neighbourhood of the expected bifurcation boundary, the present study begins with a set of uniformly distributed parameters covering the entire two-dimensional parameter domain. For each parameter combination, the corresponding solution is obtained starting from a quiescent base flow. A classification test is then applied to the computed flow field to determine whether it represents the base (symmetric) flow or whether the flow has undergone a bifurcation. The result is a labelled dataset
@@
where $\ell_{i}\in\{0,1\}$ denotes the label associated with the $i$th parameter combination, indicating nonbifurcated ($\ell_{i}=0$) or bifurcated ($\ell_{i}=1$) behaviour respectively. The labelling procedure follows exactly that described in [18] and is not repeated here. This initial dataset serves as the input to the classifier network described in Section 3.2, which is subsequently refined through adaptive sampling using the generative model introduced in Section 3.3.
@@ 3.2 Classifier Network for Bifurcation Prediction
The second stage of the methodology corresponds to the generation of a surrogate model for the bifurcation structure using a neural network classifier. The network architecture follow exactly those described in [18].
@@
where $\boldsymbol{\lambda}_{\min}$ and $\boldsymbol{\lambda}_{\max}$ denote the lower and upper bounds of the parameter space, respectively. Unlike the normalization strategy adopted in [18], where each parameter component was standardized by subtracting its mean and dividing by its standard deviation, the present work employs a fixed affine scaling to the interval $1,1]^{2}$. This approach ensures consistency with the domain required by the generative model and simplifies uniform numerical treatment of all parameters. The normalized dataset is expressed as
@@ 3.3 Adaptive Sampling with Flow-Based Deep Generative Model
The second component of the proposed approach is a flow-based deep generative model, referred to as KRnet [20, 21]. KRnet is a type of normalizing flow that learns an explicit probability density function (PDF) by constructing an invertible mapping between a simple latent space and a complex target space. This invertibility enables both the evaluation of sample likelihoods and the efficient generation of new samples. Compared with adversarial or variational approaches [8, 7, 10], normalizing flow models provide a tractable likelihood formulation, allowing their parameters to be optimized directly through likelihood maximization. However, since the parameter space in the present framework is bounded, the probability distribution to be modeled is supported on a compact domain, whereas flow-based generative models are generally defined over the entire space. To handle the bounded nature of the parameter domain, a smooth bijective mapping is introduced to transform it into an unbounded space where the flow transformation is applied.
@@
Here, the mapping $f$ is realized using the KRnet architecture, which employs a triangular transformation based on the Knothe–Rosenblatt rearrangement [1]. This design makes KRnet easily scalable to higher-dimensional sampling problems, which is particularly advantageous when the parameter space extends beyond two dimensions.
@@ 3.4 Adaptive Sampling Procedure
For completeness, we briefly summarize the initialization of both models, as this determines how their parameters evolve across adaptive stages. The classifier network is initialized only once at the beginning of the adaptive procedure: all hidden-layer weights and biases are drawn from a zero-mean normal distribution, while the output-layer weights and biases are drawn from a uniform distribution on [1,1][-1,1]. KRnet is likewise initialized only once. The neural networks appearing in the affine coupling layers use Glorot normal initialization [6] for their weights, and zero initialization for their biases. Once this initialization is performed, the parameters obtained at each stage are retained and used as the starting point for the subsequent adaptive stage.
@@ 4.1 Symmetry-Breaking Channel Flow
For this problem, the initial dataset is generated on a uniform grid in the $(\mathcal{R},\delta)$ parameter space, consisting of $8$ uniformly spaced Reynolds numbers in the range $\mathcal{R}\in[160,230]$ and $16$ logarithmically spaced inlet perturbations $\delta\in[10^{-5},10^{-2}]$, providing $128$ labeled samples. These labels are obtained from Navier–Stokes solutions computed using the IFISS software and classified according to the symmetry criterion of [18]. The classifier neural network used in this example has a single hidden layer with $32$ neurons and a softmax output layer, trained for $n_{e}^{(c)}=3000$ epochs using stochastic gradient descent (batch size $1$, learning rate $0.15$). To construct the initial training set for KRnet, a set of $n_{\text{cand}}=6000$ additional parameter points sampled uniformly in the normalized domain is evaluated by the classifier to compute the entropy weights $w(\boldsymbol{x})$. KRnet employs $6$ affine coupling layers, each equipped with two fully connected subnetworks of $24$ neurons per layer, and is trained using the WNLL loss with the ADAM optimizer (batch size $500$, learning rate $10^{-3}$) for $n_{e}^{(g)}=1000$ epochs. In total, two adaptive refinement steps are carried out, and at each step the KRnet contributes $n_{\text{new}}=50$ additional parameter points to the training set.
@@
A closer inspection of the classification patterns reveals how the adaptive strategy responds to evolving regions of uncertainty. In the initial training dataset, the classifier exhibits its greatest ambiguity near $\mathcal{R}\approx 218$ and $\log_{10}(\delta)\in[-5,-3.5]$, where the bifurcation curve cannot be reliably captured by the coarse grid alone. Consequently, KRnet focuses new samples in this neighbourhood, generating points precisely along the transition curve and producing a sharper and more clear boundary after the first adaptive step. Next, we see the uncertainty towards larger perturbations, specifically for $\log_{10}(\delta)\in[-2.75,-2.5]$ and Reynolds numbers $\mathcal{R}\approx 170$$190$. The second adaptive step again directs sampling to this newly uncertain region. These focused refinements collectively highlight the effectiveness of the adaptive sampling approach.
@@
Figure 2 shows the KRnet-generated points at the first two adaptive refinement steps. As discussed thoroughly in the methodology section, KRnet naturally concentrates samples near regions of high classifier uncertainty, which here correspond to the bifurcation boundary. The left plot shows the initial training set consisting of 128 uniformly spaced samples in the $(\mathcal{R},\log_{10}(\delta))$ parameter domain. The middle plot shows the first adaptive update, where KRnet contributes 50 new points, all lying strictly inside the admissible domain. The right plot shows the second adaptive update, containing another 50 KRnet-generated points, all of which remain strictly within the domain. Since all points consistently lie within the admissible range, no projection or filtering step was required.